The table below shows 2018 national income figures of a hypothetical economy. All figures in the table are in billion Ghana cedis. Use the information in the table to answer the questions that follows:
Employees compensation 100
Net interest. 85.2
Indirect taxes. 40.1
Subsidies. 1.2
Rental income 20
Corporate profits. 42
Fringe benefits. 3.5
Transfer payments. 0.4
Depreciation. 4.2
Net property income from abroad. -5.6
Direct taxes. 25.6
Social security contribution 12.1
a) calculate Net income, Net national product, GNP, PI, GDP
b) suppose in 2018, the population and price index of the hypothetical economy is 30,000,000 and 202, calculate GDP at constant prices and per capita income of the hypothetical economy.
1
Expert's answer
2020-08-11T08:40:56-0400

Calculating national income by income method


NDPfc= profit + rent and royalty sign + interest + mixed income + COE

= 42+20+85.2+100

= 247.2


GDPfc = NDPfc + depriciation

=. 247.2 +4.2

= 251.4

NNPfc = NDPfc + Net factor income from abroad

= 247.2+ (-5.6)

=. 241.6

GNPfc. =GDPfc + Net factor income from abroad

= 251.4+(-5.6)

=. 245.8


PI = national income + transfer payments- corporate retained earnings- income taxes - social security taxes


PI= 254.8+0.4 -(+25.6 +12.1)

= 208.5

b) per capita income= GDP/ population

= 2514000000/30000000

= 83.8


GDP at constant prices = GDP/ price index

= 251.4/202

= 1.24455446